hello, i have this question that says: the length of a rectangle is 5cm more than its width. If the width of the rectangle is y, find an expression for the area of the rectangle. How do i figure it out!? Sorry i also have another question, how do you do this one: A rhombus has side length 8x cm. One diagonal measures 12cm and the other measures 6x - 8cm. Find the area of the rhombus if area = perimeter. Thats the question and i don't have a clue how to complete it. Thankyou!
Answer by John10(297)

1)
The length of a rectangle is 5cm more than its width. If the width of the rectangle is y, find an expression for the area of the rectangle.
So the length will be y + 5 and the width is y
The area of the rectangle is A = length * width
You can finish from there:)
2)
A rhombus has side length 8x cm. One diagonal measures 12cm and the other measures 6x - 8cm. Find the area of the rhombus if area = perimeter.
Area = 12(6x - 8)/2 = 6(6x -8) = 36x - 48 (cm)^2
Perimeter = 4(8) = 32 cm
Since Area = perimeter,
36x - 48 = 32
36x = 80
x = 80/36
x = 2.2
Area = 36(2.2) - 48 = 31.2 cm^2
John10:)